# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ Abstract.
Denote by $\mathcal{P}(S)$ the semigroup obtained by equipping the family of all non-empty subsets of a (multiplicatively written) semigroup $S$ with the operation of setwise multiplication induced by $S$ itself. We call a subsemigroup $P$ of $\mathcal{P}(S)$ downward complete if any element of $S$ lies in at least one set $X\in P$ and any non-empty subset of a set in $P$ is again in $P$.
@@
We prove, for a commutative semigroup $S$, a characterization of the cancellative elements of a downward complete subsemigroup of $\mathcal{P}(S)$ in terms of the cancellative elements of $S$. Consequently, we show that, if $H$ and $K$ are cancellative semigroups and either of them is commutative, then every isomorphism from a downward complete subsemigroup of $\mathcal{P}(H)$ to a downward complete subsemigroup of $\mathcal{P}(K)$ restricts to an isomorphism from $H$ to $K$. This solves an open case of a problem of Tamura and Shafer from the late 1960s and generalizes a recent result by Bienvenu and Geroldinger, where it is assumed, among other conditions, that $H$ and $K$ are numerical monoids.
Key words and phrases:
@@
Isomorphism problems, numerical monoids, power algebras, cancellative semigroups, setwise sums and products.
2020 Mathematics Subject Classification:
Primary 20M10. Secondary 11B30, 20M13, 20M35.
1. Introduction
@@
Let $S$ be a semigroup (see the end of the section for notation and terminology). Equipped with the binary operation of setwise multiplication induced by $S$ and defined by
$(X,Y)\mapsto\{xy\colon x\in X,\kern 1.00006pty\in Y\},$
@@
the non-empty subsets of $S$ form a semigroup, herein denoted by $\mathcal{P}(S)$ and called the large power semigroup of $S$. Furthermore, the non-empty finite subsets of $S$ make up a subsemigroup of $\mathcal{P}(S)$, which we denote by $\mathcal{P}_{\mathrm{fin}}(S)$ and call the finitary power semigroup of $S$ (see Example 1.1(1) for a generalization of the same idea). Below, we will refer to either of these structures as a power semigroup of $S$.
@@
Power semigroups offer a natural algebraic framework for many important problems in additive combinatorics and related fields, from Ostmann’s conjecture [18, p. 13] on the “asymptotic additive irreducibility” of the set $\mathbb{P}$ of (positive rational) primes to Sárközy’s conjecture [21, Conjecture 1.6] on the “additive irreducibility” of the set of quadratic residues modulo $p$ for all large $p\in\mathbb{P}$. More broadly, the arithmetic of power semigroups — with a focus on questions related to the possibility or impossibility of factoring certain sets as a product of other sets that are, in a way, “irreducible” — is a rich topic in itself [9, 2, 3, 29] and has been serving as a guiding example in the development of a “unifying theory of factorization” [27, 5, 6, 7, 28] whose primary aim is to extend the classical theory [11, 12] beyond its current boundaries (so as to make it usable in non-traditional settings).
@@
As it turns out, $\mathcal{P}(S)$ was first systematically studied by Tamura and Shafer [23, 13] in the 1960s (its definition can at least be traced back to the early literature on power algebras [4, Sect. 2]) and gained substantial attention in the 1980s and 1990s due to its role in the theory of automata and formal languages [1, 20]. This led to the following question, arisen from Tamura and Shafer’s work and commonly known as the isomorphism problem for power semigroups:
Question 1.
@@
It is easy to check (see Remark 2.4(2) and cf. [29, Remark 1.1]) that, if $f\colon H\to K$ is a semigroup isomorphism, then the same is true of the function
$\mathcal{P}(H)\to\mathcal{P}(K)\colon\allowbreak X\mapsto\{f(x)\colon x\in X\}.$
Therefore, the interesting aspect of Question 2 lies entirely in the “only if” condition. E.g., the answer to the question is positive for groups [22], completely 00-simple semigroups [24, 26], and Clifford semigroups [10]; is negative for involution semigroups [8]; and appears to be unknown for cancellative semigroups, no matter whether commutative or not (cf. https://mathoverflow.net/questions/456604/).
@@
Examples 1.1.
(1) $\mathcal{P}(S)$and $\mathcal{P}_{\mathrm{fin}}(S)$ are downward complete subsemigroups of $\mathcal{P}(S)$ for any semigroup $S$; and so is the set of all one-element subsets of $S$, which is, by the way, isomorphic to $S$ through the embedding $S\to\mathcal{P}(S)\colon x\mapsto\{x\}$. More generally, let $\kappa$ be an infinite cardinal number. By basic principles of set theory (in any of the standard foundations of mathematics), each of the families $\mathcal{P}_{<\kappa}(S):=\{X\subseteq S\colon 1\leq|X|<\kappa\}\qquad\text{and}\qquad\mathcal{P}_{\leq\kappa}(S):=\{X\subseteq X\colon 1\leq|X|\leq\kappa\}$ is a downward complete subsemigroup of $\mathcal{P}(S)$. In particular, we have that $\mathcal{P}_{<\kappa}(S)=\mathcal{P}_{\mathrm{fin}}(S)$ when $\kappa$ is the cardinality of $\mathbb{N}$, and $\mathcal{P}_{\leq\kappa}(S)=\mathcal{P}(S)$ when $\kappa$ is the maximum between $|\mathbb{N}|$ and $|S|$.
(2) Let $S$ be a semigroup and $\sim$ be an equivalence relation on (the carrier set of) $S$ with the additional property that $x_{1}\sim y_{1}$ and $x_{2}\sim y_{2}$ imply $x_{1}x_{2}\sim y_{1}y_{2}$, i.e., $\sim$ is a (semigroup) congruence on $S$. Denote by $\mathcal{P}_{\sim}(S)$ the family of all non-empty sets $X\subseteq S$ such that $x\sim y$ for any $x,y\in X$.
@@
(3) Given a semigroup $S$, the intersection of any indexed family of downward complete subsemigroups of $\mathcal{P}(S)$ is itself a downward complete subsemigroup of $\mathcal{P}(S)$. The proof is straightforward, on account of the fact that a downward complete subsemigroup of $\mathcal{P}(S)$ contains, a fortiori, every one-element subset of $S$ (further details are left to the reader).
That being said, we obtain, for a commutative semigroup $S$, a characterization of the cancellative elements of a downward complete subsemigroup of $\mathcal{P}(S)$ in terms of the the cancellative elements of $S$ (Proposition 2.2); and we use it in Theorem 2.5 to prove that, if $H$ and $K$ are cancellative semigroups and either of them is commutative, then any isomorphism from a downward complete subsemigroup of $\mathcal{P}(H)$ to a downward complete subsemigroup of $\mathcal{P}(K)$ maps singletons to singletons and hence restricts to an isomorphism from $H$ to $K$ (up to the embedding mentioned in Example 1.1(1)). It follows (Corollary 2.6) that $H$ and $K$ are globally isomorphic if and only if they are isomorphic, if and only if $\mathcal{P}_{\mathrm{fin}}(H)$ is isomorphic to $\mathcal{P}_{\mathrm{fin}}(K)$. We can thus give a positive answer to Question 2 for the class of cancellative commutative semigroups and extend a recent result of Bienvenu and Geroldinger [3, Theorem 3.2(3)], where $H$ and $K$ are assumed to be numerical monoids (i.e., submonoids of the additive monoid of non-negative integers with finite complement in $\mathbb{N}$) with $\mathcal{P}_{\mathrm{fin}}(H)$ isomorphic to $\mathcal{P}_{\mathrm{fin}}(K)$ (see the comments at the end of Sect. 2).
@@ 2. Results
Let $u$ be a cancellative element in a semigroup $S$, and assume $uX=uY$ (resp., $Xu=Yu$) for some $X,Y\in\allowbreak\mathcal{P}(S)$. Given $x\in X$, there then exists $y\in Y$ such that $ux=uy$ (resp., $xu=\allowbreak yu$). It follows (by the cancellativity of $u$ in $S$) that $x=\allowbreak y$ and hence $X\subseteq\allowbreak Y$. By symmetry, we can thus conclude that $X=Y$, which ultimately shows that $\{u\}$ is a cancellative element in $\mathcal{P}(S)$.
The key idea, simple as it may be, is to establish that, in a downward complete subsemigroup of the power semigroup of a cancellative commutative semigroup $S$, the cancellative elements are all and only the one-element subsets of $S$. In fact, we will prove something slightly more general.
@@
It remains to check that every cancellative element of $P$ is a singleton. For, fix $A\in\allowbreak P$ with $|A|\geq\allowbreak 2$. We have to prove that the function $f\colon P\to P\colon X\mapsto AX$ is non-injective. We distinguish two cases.
Case 1: There exist $a,b\in A$ with $a^{2}=ab$ and $a\neq b$. Set $B:=A\smallsetminus\{a\}$. If $a\neq x\in A$, then $x\in B$ and hence $xa=ax\in AB$ (by the commutativity of $S$). Moreover, $a^{2}=ab=\allowbreak ba\in AB$ (since $a\neq b$ yields $a\in B$). It follows that $Aa\subseteq AB$ and hence $A^{2}=AB\cup Aa\subseteq AB\subseteq\allowbreak A^{2}$, namely, $f(A)=f(B)$. But this means that $f$ is not an injection (as wished), because $B$ is a non-empty proper subset of $A$ and $P$ is a downward complete subsemigroup of $\mathcal{P}(S)$ (with the result that $A,B\in P$ and $A\neq B$).
Case 2: We are not in Case 1. Pick $a,b\in A$ with $a\neq b$ (recall that $|A|\geq 2$) and set $B:=A^{2}\smallsetminus\allowbreak\{ab\}$. Since $b^{2}\neq ab$ (or else we would fall back to Case 1), we have $b^{2}\in B$ and hence $\varnothing\neq B\subsetneq A^{2}$. Given that $P$ is a downward complete subsemigroup of $\mathcal{P}(S)$, it follows that $A^{2},B\in P$ and $A^{2}\neq B$.
Now, let $x\in A$. If $xa=ab$, then it is obvious that $xab=ab^{2}\in AB$ (by the fact that $ab\neq\allowbreak b^{2}\in\allowbreak B$). Otherwise, $xa\in B$ and again $xab=bxa\in AB$ (by the commutativity of $S$). It follows that $Aab\subseteq AB$ and hence $A^{3}=AB\cup Aab\subseteq AB\subseteq A^{3}$, namely, $f(A^{2})=f(B)$. So, $A$ is a non-cancellative element of $P$ (as we have already observed that $A^{2},B\in P$ and $A^{2}\neq B$). ∎
The next example demonstrates that Proposition 2.2 does not hold for non-commutative cancellative semigroups, which suggests that addressing Question 2 in the cancellative case (assuming the answer is still affirmative) may necessitate a somewhat different approach than the one employed in this work for the cancellative commutative case.
@@
For the claim, assume $X\ast Y_{1}=X\ast Y_{2}$ for some non-empty sets $Y_{1},Y_{2}\subseteq\mathscr{F}^{+}(V)$ (the case when $Y_{1}\ast\allowbreak X=\allowbreak Y_{2}\ast X$ is symmetric and we leave it to the reader). We need to check that $Y_{1}=Y_{2}$. To this end, it is clear that, since the elements of $\mathscr{F}^{\ast}(V)$ can be uniquely represented as a (non-empty, finite, ordered) concatenation of elements of $V$, the sets $x_{1}\ast Y_{1}$ and $x_{2}\ast Y_{2}$ are disjoint for all $x_{1},x_{2}\in X$ with $x_{1}\neq x_{2}$. It follows that $X\ast Y_{1}=X\ast Y_{2}$ if and only if $x\ast Y_{1}=x\ast Y_{2}$ for each $x\in\allowbreak X$, which, by Remark 2.1 and the non-emptiness of $X$, implies $Y_{1}=Y_{2}$ (as wished).
We continue with a few additional remarks that will prove useful shortly afterwards.
Remarks 2.4.
(1) Let $f\colon H\to K$ be a semigroup isomorphism and $x$ be a left cancellative element of $H$, and pick $u,v\in K$. Since $f$ is onto, there exist $y,z\in H$ such that $u=f(y)$ and $v=f(z)$. Consequently, $f(x)u=f(x)v$ if and only if $f(xy)=f(xz)$, which implies, by the injectivity of $f$, that $xy=xz$. It is thus obvious from the left cancellativity of $x$ in $H$ that $y=z$ and hence $u=f(y)=f(z)=v$.
@@
(2) Let $H$ and $K$ be semigroups, and let $f$ be an isomorphism from a downward complete subsemigroup $P$ of $\mathcal{P}(H)$ to a downward complete subsemigroup $Q$ of $\mathcal{P}(K)$. In addition, assume $H$ is commutative. Given $u,v\in K$, there then exist $X,Y\in P$ with $f(X)=\{u\}$ and $f(Y)=\{v\}$, since $f$ is surjective and $Q$ contains every one-element subset of $K$. On the other hand, it is straightforward from the commutativity of $H$ that $XY=YX$. As a result, we have $\{u\}\{v\}=f(XY)=f(YX)=\{v\}\{u\}$ and hence $uv=vu$, which ultimately proves that $K$ is itself a commutative semigroup (cf. Proposition 1.3(a) in [19]).
(3) Assume $f\colon H\to K$ is a semigroup homomorphism and let $F$ be the function $\mathcal{P}(H)\to\mathcal{P}(K)\colon X\mapsto f[X]$, where $f[X]:=\{f(x)\colon x\in X\}\subseteq K$ is the (direct) image of $X$ under $f$. It is immediate that $F(XY)=\{f(xy)\colon x\in X,\kern 1.00006pty\in Y\}=\{f(x)f(y)\colon x\in X,\kern 1.00006pty\in Y\}=F(X)F(Y),$ for all $X,Y\in\mathcal{P}(H)$. Consequently, $F$ is a semigroup homomorphism $\mathcal{P}(H)\to\mathcal{P}(K)$.
On the other hand, if $f$ is injective and $F(X_{1})=F(X_{2})$ for some $X_{1},X_{2}\in\mathcal{P}(H)$, then $X_{1}=X_{2}$ (and hence $F$ is itself injective); otherwise, there would exist either $x\in X_{1}$ with $f(x)\in F(X_{1})\smallsetminus F(X_{2})$, or $x\in X_{2}$ with $f(x)\in F(X_{2})\smallsetminus F(X_{1})$. Moreover, if $f$ is surjective and $Y$ is a non-empty subset of $K$, then $X:=\{x\in X\colon f(x)\in Y\}$ is a non-empty subset of $H$ with $F(X)=Y$ (so that also $F$ is surjective).
Now, suppose that $f$ is an isomorphism. It then follows from the above that $F$ is an isomorphism from $\mathcal{P}(H)$ to $\mathcal{P}(K)$. In addition, $f$ being a bijection implies that a subset of $H$ and its image under $f$ have the same cardinality. So, $F$ restricts to both an isomorphism from $\mathcal{P}_{<\kappa}(H)$ to $\mathcal{P}_{<\kappa}(K)$ and an isomorphism from $\mathcal{P}_{\leq\kappa}(H)$ to $\mathcal{P}_{\leq\kappa}(K)$ for any infinite cardinal number $\kappa$ (see Example 1.1(1) for the notation). Most notably, $\mathcal{P}_{\mathrm{fin}}(H)$ is isomorphic to $\mathcal{P}_{\mathrm{fin}}(K)$.
We are finally ready to prove our main result and show, in the subsequent corollary, that Question 2 has a positive answer in the class of cancellative commutative semigroups.
@@
Proof.
Let $P$ be a downward complete subsemigroup of $\mathcal{P}(H)$ and $Q$ be a downward complete subsemigroup of $\mathcal{P}(K)$, and suppose that there is an isomorphism $f$ from $P$ to $Q$. Since either of $H$ and $K$ is commutative (by hypothesis), we have from Remark 2.4(2) that both $P$ and $Q$ are. On the other hand, each of $H$ and $K$ is cancellative. Therefore, we gather from Proposition 2.2 that the cancellative elements of $\mathcal{P}(H)$ (resp., of $\mathcal{P}(K)$) are all and only the one-element subsets of $H$ (resp., of $K$). It then follows by Remark 2.4(1) that, for every $x\in H$, there is a uniquely determined $y\in K$ such that $f(\{x\})=\allowbreak\{y\}$. Considering that the same argument also applies to the functional inverse of $f$, this suffices to conclude that $f$ restricts to an isomorphism from $H$ to $K$ (as wished). ∎
@@
Proof.
We noted in Example 1.1(1) that $\mathcal{P}(S)$ and $\mathcal{P}_{\mathrm{fin}}(S)$ are both downward complete subsemigroups of $\mathcal{P}(S)$ for every semigroup $S$. So, it is immediate from Theorem 2.5 that (a) $\Rightarrow$ (c) and (b) $\Rightarrow$ (c). This concludes the proof, as we know from Remark 2.4(3) that (c) $\Rightarrow$ (a) and (c) $\Rightarrow$ (b). ∎
@@
In fact, Bienvenu and Geroldinger’s result is actually about monoid isomorphisms (note that the finitary power semigroup of a monoid $S$ with identity $1_{S}$ is a monoid with identity $\{1_{S}\}$). However, two numerical monoids are isomorphic if and only if they are equal [14, Theorem 3]; and it is an easy to show that a surjective semigroup homomorphism $f$ from a monoid $H$ to a monoid $K$ is, a fortiori, a monoid homomorphism, namely, it maps the identity $1_{H}$ of $H$ to the identity $1_{K}$ of $K$. (Set $e:=f(1_{H})$. Each $y\in\allowbreak K$ is the image under $f$ of an element $x\in\allowbreak H$, which implies $ye=\allowbreak f(x)f(1_{H})=\allowbreak f(x1_{H})=\allowbreak f(x)=\allowbreak y$. It follows, by taking $y=\allowbreak 1_{K}$, that $f(1_{H})=\allowbreak e=\allowbreak e1_{K}=\allowbreak 1_{K}$.)

### v2

# Shijiazhuang, Hebei province, 050024 China

@@ Abstract
Let $\mathcal{P}(S)$ be the semigroup obtained by equipping the family of all non-empty subsets of a (multiplicatively written) semigroup $S$ with the operation of setwise multiplication induced by $S$ itself. We call a subsemigroup $P$ of $\mathcal{P}(S)$ downward complete if any element of $S$ lies in at least one set $X\in P$ and any non-empty subset of a set in $P$ is still in $P$.
@@
We obtain, for a commutative semigroup $S$, a characterization of the cancellative elements of a downward complete subsemigroup of $\mathcal{P}(S)$ in terms of the cancellative elements of $S$. Consequently, we show that, if $H$ and $K$ are cancellative semigroups and either of them is commutative, then every isomorphism from a downward complete subsemigroup of $\mathcal{P}(H)$ to a downward complete subsemigroup of $\mathcal{P}(K)$ restricts to an isomorphism from $H$ to $K$. This solves a special case of a problem of Tamura and Shafer from the late 1960s and generalizes a recent result by Bienvenu and Geroldinger, where it is assumed, among other conditions, that $H$ and $K$ are numerical monoids.
keywords:
@@
Isomorphism problems, numerical monoids, power algebras, cancellative semigroups, sumsets, setwise products.
1 Introduction
@@
Let $S$ be a semigroup (see the end of the section for notation and terminology). Equipped with the binary operation of setwise multiplication induced by $S$ on its own power set and defined by
$(X,Y)\mapsto\{xy\colon x\in X,\,y\in Y\},$
@@
the non-empty subsets of $S$ form a semigroup, herein denoted by $\mathcal{P}(S)$ and called the large power semigroup of $S$. Furthermore, the non-empty finite subsets of $S$ make up a subsemigroup of $\mathcal{P}(S)$, which we denote by $\mathcal{P}_{\mathrm{fin}}(S)$ and call the finitary power semigroup of $S$ (see Example 1.3 for a generalization). Below, we will refer to either of these structures as a power semigroup of $S$.
@@
Power semigroups offer a natural algebraic framework for many important problems in additive combinatorics and related fields, from Ostmann’s conjecture [18, p. 13] on the “asymptotic additive irreducibility” of the set $\mathbb{P}$ of (positive rational) primes to Sárközy’s conjecture [21, Conjecture 1.6] on the “additive irreducibility” of the set of quadratic residues modulo $p$ for all large $p\in\mathbb{P}$. More broadly, the arithmetic of power semigroups — with a focus on questions related to the possibility or impossibility of factoring certain sets as a product of other sets that are, in a way, “irreducible” — is a rich topic in itself [9] [2] [3] [29] and has been serving as a guiding example in the development of a “unifying theory of factorization” [27] [5] [6] [7] [28] whose primary aim is to extend the classical theory [11] [12] beyond its current boundaries and make it usable in non-traditional settings.
@@
As it turns out, $\mathcal{P}(S)$ was first systematically studied by Tamura and Shafer [23] [13] in the 1960s (its definition can at least be traced back to the early literature on power algebras [4, Sect. 2]) and gained substantial attention in the 1980s and 1990s due to its role in the theory of automata and formal languages [1] [20]. This led to the following question, arisen from Tamura and Shafer’s work and commonly known as the isomorphism problem for power semigroups.
Question 1.1.
@@
It is easy to check (see Remark 2.6 and cf. [29, Remark 1.1]) that, if $f\colon H\to K$ is a semigroup isomorphism, then the same is true of the function
$\mathcal{P}(H)\to\mathcal{P}(K)\colon\allowbreak X\mapsto f[X],$
where $f[X]:=\{f(x)\colon x\in X\}\subseteq K$ is the (direct) image of $X$ under $f$. Therefore, the interesting aspect of Question 1.2 lies entirely in the “only if” condition. E.g., the answer to the question is positive for groups [22], completely 00-simple semigroups [24] [26], and Clifford semigroups [10]; is negative for involution semigroups [8]; and appears to be unknown for cancellative semigroups (see https://mathoverflow.net/questions/456604/ for further discussion).
@@
Example 1.3.
$\mathcal{P}(S)$ and $\mathcal{P}_{\mathrm{fin}}(S)$ are downward complete subsemigroups of $\mathcal{P}(S)$ for any semigroup $S$; and so is the set of all one-element subsets of $S$, which is obviously isomorphic to $S$ through the embedding $S\to\mathcal{P}(S)\colon x\mapsto\{x\}$.
More generally, let $\kappa$ be an infinite cardinal number. By basic principles of set theory (in any of the standard foundations of mathematics), each of the families
$\mathcal{P}_{<\kappa}(S):=\{X\subseteq S\colon 1\leq|X|<\kappa\}$
and
$\mathcal{P}_{\leq\kappa}(S):=\{X\subseteq S\colon 1\leq|X|\leq\kappa\}$
is a downward complete subsemigroup of $\mathcal{P}(S)$. In particular, we have that $\mathcal{P}_{<\kappa}(S)=\mathcal{P}_{\mathrm{fin}}(S)$ when $\kappa$ is the cardinality of $\mathbb{N}$, and $\mathcal{P}_{\leq\kappa}(S)=\mathcal{P}(S)$ when $\kappa$ is the maximum between $|\mathbb{N}|$ and $|S|$.
Example 1.4.
Let $S$ be a semigroup and $\sim$ be an equivalence relation on (the carrier set of) $S$ with the additional property that $x_{1}\sim y_{1}$ and $x_{2}\sim y_{2}$ imply $x_{1}x_{2}\sim y_{1}y_{2}$, i.e., $\sim$ is a (semigroup) congruence on $S$. Denote by $\mathcal{P}_{\sim}(S)$ the family of all non-empty sets $X\subseteq S$ such that $x\sim y$ for any $x,y\in X$.
@@
Example 1.5.
Given a semigroup $S$, the intersection of any indexed family of downward complete subsemigroups of $\mathcal{P}(S)$ is itself a downward complete subsemigroup of $\mathcal{P}(S)$. The proof is straightforward, on account of the fact that a downward complete subsemigroup of $\mathcal{P}(S)$ contains, a fortiori, every one-element subset of $S$ (further details are left to the reader).
That being said, we obtain, for a commutative semigroup $S$, a characterization of the cancellative elements of a downward complete subsemigroup of $\mathcal{P}(S)$ in terms of the the cancellative elements of $S$ (Proposition 2.2); and we use it in Theorem 2.8 to prove that, if $H$ and $K$ are cancellative semigroups and either of them is commutative, then any isomorphism from a downward complete subsemigroup of $\mathcal{P}(H)$ to a downward complete subsemigroup of $\mathcal{P}(K)$ maps singletons to singletons and hence restricts to an isomorphism from $H$ to $K$ (up to the embedding mentioned in Example 1.3). It follows (Corollary 2.10) that $H$ and $K$ are globally isomorphic if and only if they are isomorphic, if and only if $\mathcal{P}_{\mathrm{fin}}(H)$ is isomorphic to $\mathcal{P}_{\mathrm{fin}}(K)$. We can thus give a positive answer to Question 1.2 for the class of cancellative commutative semigroups and extend a recent result of Bienvenu and Geroldinger [3, Theorem 3.2(3)], where $H$ and $K$ are numerical monoids (i.e., submonoids of the additive monoid of non-negative integers with finite complement in $\mathbb{N}$) with $\mathcal{P}_{\mathrm{fin}}(H)$ isomorphic to $\mathcal{P}_{\mathrm{fin}}(K)$ (see the comments at the end of Sect. 2).
@@ 2 Results
Let $u$ be a left (resp., right) cancellative element in a semigroup $S$, and assume $uX=uY$ (resp., $Xu=Yu$) for some $X,Y\in\allowbreak\mathcal{P}(S)$. Given $x\in X$, there then exists $y\in Y$ such that $ux=uy$ (resp., $xu=\allowbreak yu$). It follows (by the hypothesis on $u$) that $x=\allowbreak y$ and hence $X\subseteq\allowbreak Y$. By symmetry, we can thus conclude that $X=Y$, which ultimately shows that $\{u\}$ is a left (resp., right) cancellative element in $\mathcal{P}(S)$.
The key idea, simple as it may be, is to establish that, in a downward complete subsemigroup of the large power semigroup of a cancellative commutative semigroup $S$, the cancellative elements are all and only the one-element subsets of $S$. In fact, we will prove something slightly more general.
@@
It remains to check that every cancellative element of $P$ is a singleton. For, fix $A\in\allowbreak P$ with $|A|\geq\allowbreak 2$. We have to prove that the function
$f\colon P\to P\colon X\mapsto AX$
is non-injective. We distinguish two cases.
Case 1: There exist $a,b\in A$ with $a^{2}=ab$ and $a\neq b$. Set $B:=A\setminus\{a\}$. If $a\neq x\in A$, then $x\in B$ and hence $xa=ax\in AB$ (by the commutativity of $S$). Moreover, $a^{2}=ab\in AB$ (since $a\neq b$ yields $b\in B$). It follows that $Aa\subseteq AB$ and hence $A^{2}=AB\cup Aa\subseteq AB\subseteq\allowbreak A^{2}$, namely, $f(A)=f(B)$. But this means that $f$ is not an injection (as wished), because $P$ is a downward complete subsemigroup of $\mathcal{P}(S)$ and $B$ is a non-empty proper subset of $A$ (with the result that $A,B\in P$ and $A\neq B$).
Case 2: We are not in Case 1. Pick $a,b\in A$ with $a\neq b$ (recall that $|A|\geq 2$) and set $B:=A^{2}\setminus\allowbreak\{ab\}$. Since $b^{2}\neq ab$ (or else we would fall back to Case 1), we have $b^{2}\in B$ and hence $\emptyset\neq B\subsetneq A^{2}$. Given that $P$ is a downward complete subsemigroup of $\mathcal{P}(S)$, it follows that $A^{2},B\in P$ and $A^{2}\neq B$.
Now, let $x\in A$. If $xa=ab$, then it is obvious that $xab=ab^{2}\in AB$ (by the fact that $ab\neq\allowbreak b^{2}\in\allowbreak B$). Otherwise, $xa\in B$ and again $xab=bxa\in AB$ (by the commutativity of $S$). Consequently, we find that $Aab\subseteq AB$ and hence $A^{3}=AB\cup Aab\subseteq AB\subseteq A^{3}$, namely, $f(A^{2})=f(B)$. So, $A$ is a non-cancellative element of $P$ (as we have already observed that $A^{2},B\in P$ and $A^{2}\neq B$).
The next example demonstrates that Proposition 2.2 does not hold for non-commutative cancellative semigroups. This suggests that addressing Question 1.2 in the cancellative, non-commutative case (assuming the answer is still affirmative) may necessitate a somewhat different approach than the one envisaged in this work for the commutative case.
@@
For the claim, assume $X\ast Y_{1}=X\ast Y_{2}$ for some non-empty sets $Y_{1},Y_{2}\subseteq\mathscr{F}^{+}(V)$ (the case when $Y_{1}\ast\allowbreak X=\allowbreak Y_{2}\ast X$ is symmetric). We need to check that $Y_{1}=Y_{2}$. To this end, it is clear that, since the elements of $\mathscr{F}^{+}(V)$ can be uniquely represented as a (non-empty, finite, ordered) concatenation of elements of $V$, the sets $x_{1}\ast Y_{1}$ and $x_{2}\ast Y_{2}$ are disjoint for all $x_{1},x_{2}\in X$ with $x_{1}\neq x_{2}$. It follows that $X\ast Y_{1}=X\ast Y_{2}$ if and only if $x\ast Y_{1}=x\ast Y_{2}$ for each $x\in\allowbreak X$, which, by Remark 2.1 and the non-emptiness of $X$, implies $Y_{1}=Y_{2}$ (as wished).
We continue with a few additional observations that will prove useful shortly afterwards.
Remark 2.5.
Let $f\colon H\to K$ be a semigroup isomorphism and $x$ be a left cancellative element of $H$, and pick $u,v\in K$. Since $f$ is onto, there exist $y,z\in H$ such that $u=f(y)$ and $v=f(z)$. Consequently, $f(x)u=f(x)v$ if and only if $f(xy)=f(xz)$, which implies, by the injectivity of $f$, that $xy=xz$. It is thus obvious from the left cancellativity of $x$ in $H$ that $y=z$ and hence $u=f(y)=f(z)=v$.
@@
Remark 2.6.
Let $H$ and $K$ be semigroups, and let $f$ be an isomorphism from a downward complete subsemigroup $P$ of $\mathcal{P}(H)$ to a downward complete subsemigroup $Q$ of $\mathcal{P}(K)$. In addition, assume $H$ is commutative. Given $u,v\in K$, there then exist $X,Y\in P$ with $f(X)=\{u\}$ and $f(Y)=\{v\}$, since $f$ is surjective and $Q$ contains every one-element subset of $K$. On the other hand, it is straightforward from the commutativity of $H$ that $XY=YX$. As a result, we have $\{u\}\{v\}=f(XY)=f(YX)=\{v\}\{u\}$ and hence $uv=vu$, which ultimately proves that $K$ is itself a commutative semigroup (cf. Proposition 1.3(a) in [19]).
Remark 2.7.
Assume $f\colon H\to K$ is a semigroup homomorphism and let $F$ be the function $\mathcal{P}(H)\to\mathcal{P}(K)\colon X\mapsto f[X]$. It is immediate that
$\begin{split}F(XY)&=\{f(xy)\colon x\in X,\,y\in Y\}=\{f(x)f(y)\colon x\in X,\,y\in Y\}=F(X)F(Y),\end{split}$
for all $X,Y\in\mathcal{P}(H)$. Consequently, $F$ is a homomorphism $\mathcal{P}(H)\to\mathcal{P}(K)$.
On the other hand, if $f$ is injective and $F(X_{1})=F(X_{2})$ for some $X_{1},X_{2}\in\mathcal{P}(H)$, then $X_{1}=X_{2}$, and hence $F$ is itself injective; otherwise, there would exist either $x\in X_{1}$ with $f(x)\in F(X_{1})\setminus F(X_{2})$ or $x\in X_{2}$ with $f(x)\in F(X_{2})\setminus F(X_{1})$. Moreover, if $f$ is surjective and $Y$ is a non-empty subset of $K$, then $X:=\{x\in X\colon f(x)\in Y\}$ is a non-empty subset of $H$ with $F(X)=Y$, which ultimately means that also $F$ is surjective.
Now, suppose that $f$ is an isomorphism. It then follows from the above that $F$ is an isomorphism from $\mathcal{P}(H)$ to $\mathcal{P}(K)$. In addition, $f$ being a bijection implies that a subset of $H$ and its image under $f$ have the same cardinality.
It is therefore clear (see Example 1.3 for the notation) that, for every infinite cardinal $\kappa$, the function $F$ restricts to an isomorphism $\mathcal{P}_{<\kappa}(H)\to\mathcal{P}_{<\kappa}(K)$ as well as to an isomorphism $\mathcal{P}_{\leq\kappa}(H)\to\mathcal{P}_{\leq\kappa}(K)$. Most notably, $F$ restricts to an isomorphism from $\mathcal{P}_{\mathrm{fin}}(H)$ to $\mathcal{P}_{\mathrm{fin}}(K)$.
We are finally ready to prove our main result and show, in the subsequent corollary, that Question 1.2 has a positive answer in the class of cancellative commutative semigroups.
@@
Proof 2.9.
Let $P$ be a downward complete subsemigroup of $\mathcal{P}(H)$ and $Q$ be a downward complete subsemigroup of $\mathcal{P}(K)$, and suppose that there is an isomorphism $f$ from $P$ to $Q$. Since either of $H$ and $K$ is commutative (by hypothesis), we have from Remark 2.6 that both $P$ and $Q$ are. On the other hand, each of $H$ and $K$ is cancellative. Therefore, we gather from Proposition 2.2 that the cancellative elements of $\mathcal{P}(H)$ (resp., of $\mathcal{P}(K)$) are all and only the one-element subsets of $H$ (resp., of $K$). It then follows by Remark 2.5 that, for every $x\in H$, there is a uniquely determined $y\in K$ such that $f(\{x\})=\allowbreak\{y\}$. Considering that the same argument also applies to the functional inverse of $f$, this suffices to conclude that $f$ restricts to an isomorphism from $H$ to $K$ (as wished).
@@
Proof 2.11.
We noted in Example 1.3 that $\mathcal{P}(S)$ and $\mathcal{P}_{\mathrm{fin}}(S)$ are both downward complete subsemigroups of $\mathcal{P}(S)$ for every semigroup $S$. So, it is immediate from Theorem 2.8 that (a) $\Rightarrow$ (c) and (b) $\Rightarrow$ (c). This concludes the proof, as we know from Remark 2.7 that (c) $\Rightarrow$ (a) and (c) $\Rightarrow$ (b).
@@
In fact, Bienvenu and Geroldinger’s result is actually about monoid isomorphisms (note that the finitary power semigroup of a monoid $S$ with identity $1_{S}$ is a monoid with identity $\{1_{S}\}$). However, two numerical monoids are isomorphic if and only if they are equal [14, Theorem 3]; and it is easy to show that a surjective semigroup homomorphism $f$ from a monoid $H$ to a monoid $K$ is, a fortiori, a monoid homomorphism, i.e., it maps the identity $1_{H}$ of $H$ to the identity $1_{K}$ of $K$.
For, set $e:=f(1_{H})$. Each $y\in\allowbreak K$ is the image under $f$ of an element $x\in\allowbreak H$, which implies $ye=\allowbreak f(x)f(1_{H})=\allowbreak f(x1_{H})=\allowbreak f(x)=\allowbreak y$. It follows, by taking $y=\allowbreak 1_{K}$, that $f(1_{H})=\allowbreak e=\allowbreak 1_{K}e=\allowbreak 1_{K}$.